YESTERDAY YOU BOUGHT A HUGE WATERMELON THAT WEIGHS 100 POUNDS. A WATERMELON IS 90% WATER. YOU LEFT IT OUTSIDE IN THE HOT SUN. SOME OF THE WATER EVAPORATED, SO IT IS NOW 80% WATER. HOW MUCH DOES THE WATERMELON WEIGH NOW?
Answer by stanbon(48568)

YESTERDAY YOU BOUGHT A HUGE WATERMELON THAT WEIGHS 100 POUNDS. A WATERMELON IS 90% WATER. YOU LEFT IT OUTSIDE IN THE HOT SUN. SOME OF THE WATER EVAPORATED, SO IT IS NOW 80% WATER. HOW MUCH DOES THE WATERMELON WEIGH NOW?
--------------
The melon starts with 0.9*100 = 90 lbs of water and 10 lbs of "other".
--------
After reduction its "other" is 20% of its weight.
----------
Let it's weight now be "x" lbs.
EQUATION:
0.20x = 10 lbs
x = 10/0.2 = 50 lbs. (it's weight now is 50 lbs.)
==================
Cheers,
Stan H.